OpenStudy (australia10):

how is 0.3 repeating 1/3 3 tenths is written as 3/10 Also 0.6 repeating is apparently 2/3 but how? 6/10=3/5?

zepdrix (zepdrix):

Let's call the number 0.3 repeating like... x or something, \(\large\rm x=0.333...\) If we multiply both sides of this assignment by 10, \(\large\rm 10x=3.333...\) The multiplication just moves the decimal one place, ya? Then let's apply subtraction,\[\large\rm 10x-x=3.333...-0.333...\]On the left side of our equation, we have 9x, on the right side, they have the same decimal portion, so those will subtract off,\[\large\rm 9x=3\]

zepdrix (zepdrix):

Then divide to get the value for x,\[\large\rm x=\frac39\]

zepdrix (zepdrix):

Which simplifies to 1/3. Kind of a neat little trick, ya? :)

OpenStudy (australia10):

Where are you getting the other x from?

OpenStudy (australia10):

Thanks for helping just not sure how you worked it out.

OpenStudy (otherworldly):

do u want to know how 6/10 = 3/5